OpenStudy (anonymous):

PLEASE HELP ME!!!! WILL FAN+MEDAL!!!! Which of the following is a graph of y=x^2

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b

OpenStudy (anonymous):

c

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Can you help with my last question? @Butterfield1215

OpenStudy (welshfella):

y = x^2 will touch the origin and open upwards

OpenStudy (welshfella):

and also pass through the points (1,1) and (-1,1)

OpenStudy (welshfella):

and also pass through (2 , 4) because when x = 2 y = 4 (4 = 2^2)

OpenStudy (anonymous):

So, would it be B? @welshfella

OpenStudy (anonymous):

@welshfella